# Performant Automatic BLAS Offloading on Unified Memory Architecture with OpenMP First-Touch Style Data Movement

### v2

@@ Abstract.
BLAS is a fundamental building block of advanced linear algebra libraries and many modern scientific computing applications. GPU is known for its strong arithmetic computing capability, and highly suited for BLAS operations. However, porting code to GPUs often requires significant effort especially for large complex codes or legacy codes, even for BLAS heavy applications. While various tools exist to automatically offload BLAS to GPU, they are often impractical due to the high costs associated with mandatory data transfers. The advent of unified memory architectures in recent GPU designs, such as the NVIDIA Grace-Hopper, allows cache-coherent memory access across all types of memory for both CPU and GPU, potentially eliminating the bottlenecks faced in conventional architectures. This breakthrough paves the way for innovative application developments and porting strategies. In this paper, building on my preliminary work(Li et al., 2024) demonstrating the possibility of performant automatic *gemm offload, I extend the framework to all level-3 BLAS operations, and present SCILIB-Accel(Li and Wang, 2024), a novel tool for automatic BLAS offload . SCILIB-Accel leverages the cache-coherent NVLink C2C interconnect in Grace-Hopper and introduces a Device First-Use data movement policy. This policy, inspired by the OpenMP First-Touch approach in multi-socket CPU programming, minimizes CPU-GPU data transfers for typical scientific computing codes. Additionally, utilizing the dynamic binary instrumentation technique, the tool intercepts BLAS symbols directly from a CPU binary, requiring no code modifications or recompilation. SCILIB-Accel has been evaluated using multiple quantum physics codes on up to a few hundred GPU nodes, yielding promising speedups. Notably, for the LSMS method in the MuST suite, a 3x speedup was achieved on Grace-Hopper compared to Grace-Grace. SCILIB-Accel is the first tool to deliver practical, high-performance automatic BLAS offload for scientific applications.
@@
journalyear: 2024copyright: acmlicensedconference: xxxxxxxx; Dec 21–25, 2024; Austin, TX, USAbooktitle: XXXXXXXccs: General and reference Performanceccs: Mathematics of computing Mathematical software performanceccs: Software and its engineering Software performance
@@ 1. Introduction
The Basic Linear Algebra Subprograms (BLAS) serves as a building block for many scientific computing applications and forms the foundation for advanced linear algebra libraries such as LAPACK and ScaLAPACK. These libraries are extensively used in mathematical software like Mathematica and MATLAB, as well as in data science packages such as NumPy, and in computational chemistry and physics applications. Notably, BLAS is heavily utilized in quantum chemistry and quantum physics codes, as linear algebra is the natural language of quantum mechanics.
Modern general-purpose Graphics Processing Units (GPUs) are known for their exceptional arithmetic compute power. Their raw FP32 and FP64 compute capabilities significantly outpace those of CPUs, making GPUs an ideal platform for running BLAS-intensive applications. While all major GPU manufacturers provide highly optimized BLAS libraries, such as cuBLAS for NVIDIA GPUs and rocBLAS for AMD GPUs, these GPU libraries have slightly different interfaces and are not drop-in replacements for CPU BLAS libraries. More critically, using these GPU BLAS libraries, like any other GPU porting task, requires developers to manually manage data movement for optimal performance. Consequently, porting large codes, legacy codes and codes with complex workflow to GPU isn’t trivial, sometimes daunting and requires significant investment of manpower. Furthermore, supercomputers are becoming increasingly GPU-centric due to the rapid advancement of GPUs and the surge of AI applications, this trend generates a pressing needs to port more scientific codes to GPU, but many researchers lack the expertise for GPU porting and face challenges securing funding for pure code-porting efforts. Given BLAS’ central role and its extensive GPU support, there have been numerous attempts to automate GPU usage, as outlined in Section 2.2. However, limitations inherent to conventional GPU architectures often necessitate frequent data transfers between main memory and GPU memory, resulting in overheads that are unacceptable for practical use.
@@
In my previous work(Li et al., 2024; Li and Wang, 2024), I presented a proof-of-concept framework for doing symbol interception and replacement with Dynamic Binary Instrumentation (DBI), and preliminary tests of automatically offloading the *gemm (general matrix multiplication) routines, proving that performat automatic BLAS offload is achievable. In this paper, I further extend the implementation to all level-3 BLAS calls, and focus on discussing a novel data management strategy named Device First-use policy. Inspired by the OpenMP First-Touch memory management approach in multi-socket CPU or NUMA programming, the Device First-Use policy is designed for CPU-GPU systems where data is moved to GPU memory (or general device memory) upon its first use by a GPU kernel. This approach is both simple and effective, minimizing data transfers in practical BLAS use cases. Tests across several BLAS-intensive scientific applications on up to hundreds of GPU nodes demonstrate significant speedups. Although all tests are done on the NVIDIA Grace-Hopper system, the methodology is generic for any CPU-GPU system with cache coherency.
The remainder of this paper is organized as follows: Section 2 reviews the NVIDIA Grace-Hopper unified memory architectures and related BLAS offloading work. Section 3 details the implementation of SCILIB-Accel. In Section 4, I apply the tool to two BLAS-intensive scientific computing codes on up to 200 Grace-Hopper nodes and discuss the results along with performance issues of the NVIDIA Grace-Hopper system. Finally, Section 5 concludes the paper.
@@ 2.1. Coherent memory in NVIDIA Grace-Hopper
Although the two memory subsystems can be cache coherently accessed by both CPU and GPU, the bandwidth varies significantly across different access patterns, as shown in Table 1. When the CPU accesses its local LPDDR5X memory, it achieves a descent bandwidth of over 400 GB/s. The GPU accessing its local HBM3 memory delivers even more impressive performance, reaching 3.6 TB/s. The NVLink-C2C interconnect provides 450 GB/s of bandwidth in each direction, which adequately supports the full bandwidth of LPDDR5X, allowing GPU access to remote LPDDR5X memory at 400+ GB/s. However, CPU access to HBM3 memory is substantially slower, achieving only approximately 140 GB/s. These bandwidth disparities indicate that data locality remains crucial for Grace-Hopper systems. Developers must carefully optimize data movement patterns to maximize performance, similar to conventional GPU architectures, rather than relying solely on coherent unified memory access for production workloads.
@@ 3. Performant Automatic BLAS Offload
Figure 1. Workflow of Automatic BLAS Offload. The BLAS in the CPU binary is intercepted and replaced with a BLAS wrapper where GPU BLAS call is maded and data is moved between CPU resident memory and GPU resident memory.
A basic workflow of automatic BLAS offload tool is illustrated in Figure 1 where an dgemm call in the caller code is intercepted and redirected to a BLAS wrapper that manages data movement and makes the GPU BLAS call. Conceptually this workflow contains two tasks: 1) intercept BLAS symbols and replace them with a BLAS wrapper where GPU BLAS is called, 2) manage data movement between CPU and GPU resident memories. In the following content, I will discuss how these tasks are implemented in SCILIB-Accel, most critially how the data movement can be optimized using a novel data movement strategy inspired by the OpenMP First-Touch data placement policy.
@@ 3.1. Symbol Interception
For automatic BLAS offload, I intercept BLAS calls where the trampoline function (BLAS wrapper function) maintains the same signature as the original function. This results in minimal overhead, as the register data for function arguments remain undisturbed. This mechanism finds extensive use in profilers, and here I use the PEAK(Wang and Li, 2023) lightweight profiler that was developed by me and my collaborator as a DBI framework, ensuring portability across various architectures including but not limited to x86 and ARM. Inside the DBI framework, the FRIDA-GUM(Frida, nd) binary instrumentation backbend is chosen for simplicity and performance.
@@ 3.2. Data Movement Strategies
Managing data movement is often the most critical part of GPU porting as data transfer speed has been a limiting factor. It is even more so for developing an automatic offload tool, as the tool deals with pure CPU code that is totally untuned for GPU. Here, I discuss three data movement strategies that utilize different features of Grace-Hopper so that this helps us better understand the challenge and opportunities for doing automatic offload on unified memory architecture.
@@ 3.2.1. Strategy 1, Mem-Copy
void my_dgemm(…) {
cudaMalloc(&device_A,size_A);
cudaMalloc(&device_B,size_B);
cudaMalloc(&device_C,size_C);
cudaMemcpy(device_A,host_A,size_A,cudaMemcpyDefault);
cudaMemcpy(device_B,host_B,size_B,cudaMemcpyDefault);
cublasDgemm(handle, TransA, TransB, m, n, k,
&alpha, device_A, lda, device_B, ldb, &beta,
device_C, ldc);
cudaMemcpy(host_C,device_C,size_C,cudaMemcpyDefault);
@@ 3.2.2. Strategy 2, counter-based migration
Since the CPU resident memory (LPDDR5X) and the GPU resident memory (HBM3) are physically unified with cache-coherent NVLink C2C, CPU matrix pointers can be passed directly to cuBLAS calls. One can also use numactl to force all memory to be resident on the HBM, but given the poor bandwidth of CPU access HBM, a performance penalty is expect. A new feature in Grace-Hopper designed to better serve the unified memory is the access counter on Hopper GPU, namely the CUDA runtime will automatically move memory pages from LPDDR5X to HBM3 based on remote memory accesses detected by the counter. This counter-based migration mechanism can serve automatic offload when CPU resident matrices are passed to a GPU kernel. A pseudocode example is shown below.
@@
void my_dgemm(…) {
cublasDgemm(handle, TransA, TransB, m, n, k,
@@ 3.2.3. Strategy 3, Device First-Use policy
One alternative way to look at the Grace-Hopper superchip is that is operates as a heterogeneous dual-socket system, with one socket being a CPU and the other a GPU. Its NUMA configuration also mirrors that of a dual-socket CPU system, where CPU-resident memory is assigned to NUMA 0 and GPU-resident memory to NUMA 1. The data management challenge here in CPU-GPU superchip is a reminiscence of the OpenMP First-Touch data placement policy used in CPU-CPU NUMA programming. While malloc calls can theoretically be intercepted to allocate memory directly for GPU access, these calls originate from the CPU binary, making it impractical to identify which memory regions will later be used by the GPU. As a result, implementing a GPU-first-touch policy is not feasible. To address this, I propose a GPU First-Use policy, where data is moved to the GPU the first time it is accessed by a CUDA kernel. More generally, this concept can be extended to a Device First-Use policy, applicable to any accelerator device with cache-coherent memory access. Table 2 summarizes the features of Device First-Use and highlights its similarities to OpenMP First-Touch.
@@
void my_dgemm(…) {
@@
if (host-C on NUMA 0) move_pages (host_C, to NUMA 1) ;
cublasDgemm(handle, TransA, TransB, m, n, k,
@@ 3.3. Usage Instructions
    data management strategies with Device First-Use as default
@@
    minimum matrix size that will allow a BLAS call to be offloaded, if matrices are small, then the BLAS call will stay on CPU. The default threshold is $N_{avg}>500$ where $N_{avg}$ is the average matrix size, the definition of $N_{avg}$ is routine dependent. For general matrix multiplication routines C=AxB, $N_{avg}=(MNK)^{1/3}$ where dimensions of matrices A, B and C are $M\times K$, $K\times N$ and $M\times N$. The default threshold is a safe lower-bound based on preliminary dgemm testings on Grace-Hopper and can certainly be further fine-tuned for different kernels or precisions.
@@ 4. Performance Testings and Discussion
In this section, I perform application tests using production HPC codes. Since quantum chemistry/physics are known to heavily rely on BLAS operations, as linear algebra is the natural language of quantum mechanics, two codes, MuST and PARSEC, from this scientific domain are chosen. These codes are part of the Characteristic Science Application (CSA)(Foundation, 2021) efforts funded the National Scientific Foundation (NSF) as representative workloads for the Leadership Class Computing Facility.
@@
It is important to emphasize that all performance comparisons presented here are based on an equal number of nodes: Grace-Grace nodes (two CPUs with 144 cores) and Grace-Hopper nodes (one CPU and one GPU). This approach ensures both simplicity and fairness in the analysis. Power and cost are also key factors in performance studies. The power consumption of a Grace-Hopper node is approximately twice that of a Grace-Grace node under full load. Additionally, for each node-hour, the Service Unit (SU) charged to users for Grace-Hopper nodes is three times higher than that for Grace-Grace nodes. This charge rate reflects the costs associated with acquiring, maintaining, and supporting these different types of nodes.
@@ 4.4. Performance Issues with Grace-Hopper Relevant to Auto Offload
In this part of the paper, I discuss a few performance issues observed in the above application test, these issues reflect the immaturity of the software or hardware design in the current Grace-Hopper system, and could be fixed in the future to further improve performance and usability of automatic offload.
@@ 4.4.1. Counter-based page migration
Since the data migrated to GPU will not be migrated back, automatic offload with counter-based migration should work very similar to manually implemented Device First-Use policy, but we have already seen the slow performance of the counter-based migration in application tests. Here I present a few simple dgemm test cases with different matrix sizes, and provide a deeper understanding of the issues with the counter-based migration. In the following test, matrices A(M,K), B(K,N) and C(M,N) are allocated by malloc and initialized on CPU resident memory, then multiplication $C=A\times B$ is performed at least 5 times by passing these matrices to cublasDgemm, so that the access-counter should allow the matrices to be migrated to GPU resident memory. The NUMA locations of the matrices are reported after each cublasDgemm call, and runtime for each call is reported. From the NUMA location, we can infer whether the data is on the CPU resident memory (NUMA 0) or GPU resident memory (NUMA 1). Four sets of inputs with different matrix sizes and shapes are tested, all of which use FP64 precision.
@@
The case gets strange for non-square matrix dimensions. When M=32, N=2400 and K=93536, a matrix size commonly used in the PARSEC workload, the matrix size becomes 24.0MB for A, 1795.9MB for B, and 0.6MB for C. Throughout my test, only matrix A gets migrated to GPU, while B and C are always on the CPU. This is very counter-intuitive as one would expect the bigger matrix B should at least be moved, as it generates the largest volume of remote memory accesses, but NVIDIA’s counter-based migration algorithm fails to move it.
From these tests, we can see that NVIDIA’s algorithm tends to migrate smaller data to GPU. The reason could be that the algorithm only makes the decision based on a single CUDA kernel call, i.e. moving the big data is costly comparing to that single CUDA kernel runtime, so it decides not to migrate disregarding the fact that the big data gets accessed by the GPU many more times later on. The current counter-based migration is thus unpredictable and inconsistent, and NVIDIA should provide a simple way to disable it by users instead of requiring to unload a kernel model by root.

### v3

@@ Abstract.
BLAS is critical for scientific computing. While GPUs excel at BLAS operations, porting code to GPUs remains challenging due to code complexity and data transfer overhead. Unified memory architectures, like NVIDIA Grace-Hopper, enable cache-coherent memory access between CPU and GPU, addressing conventional bottlenecks. Building on our preliminary work for gemm offload, we present SCILIB-Accel, a novel tool for automatic offloading of all level-3 BLAS routines. SCILIB-Accel leverages Grace-Hopper’s NVLink C2C interconnect and introduces a Device First-Use policy, inspired by OpenMP’s First-Touch approach, to minimize CPU-GPU transfers. Using dynamic binary instrumentation, it intercepts BLAS calls at runtime without code modifications or recompilation. Evaluated on quantum physics codes across hundreds of GPU nodes, SCILIB-Accel achieves significant speedups, including a 3× improvement for the LSMS method in the MuST suite on Grace-Hopper versus Grace-Grace. SCILIB-Accel is the first tool enabling practical, high-performance automatic BLAS offload for scientific applications.
@@
journalyear: 2025copyright: acmlicensedconference: xxxxxxxx; Nov 16–21, 2025; St. louis, MO, USAbooktitle: XXXXXXXccs: General and reference Performanceccs: Mathematics of computing Mathematical software performanceccs: Software and its engineering Software performance
@@ 1. Introduction
The Basic Linear Algebra Subprograms (BLAS) serve as a building block for many scientific computing applications and form the foundation for advanced linear algebra libraries such as LAPACK and ScaLAPACK. These libraries are extensively used in mathematical software like Mathematica and MATLAB, as well as in data science packages such as NumPy, and in computational chemistry and physics applications. Notably, BLAS is heavily utilized in quantum chemistry and quantum physics codes, as linear algebra is the natural language of quantum mechanics.
Modern general-purpose Graphics Processing Units (GPUs) are known for their exceptional arithmetic compute power. Their raw FP32 and FP64 compute capabilities significantly outpace those of CPUs, making GPUs an ideal platform for running BLAS-intensive applications. While all major GPU manufacturers provide highly optimized BLAS libraries, such as cuBLAS for NVIDIA GPUs and rocBLAS for AMD GPUs, these GPU libraries have slightly different interfaces and are not drop-in replacements for CPU BLAS libraries. More critically, using these GPU BLAS libraries, like any other GPU porting task, requires developers to manually manage data movement for optimal performance. Consequently, porting large codes, legacy codes and codes with complex workflow to GPU isn’t trivial, sometimes daunting,, and requires significant investment of manpower. Furthermore, supercomputers are becoming increasingly GPU-centric due to the rapid advancement of GPUs and the surge of AI applications, this trend generates pressing needs to port more scientific codes to GPU, but many researchers lack the expertise for GPU porting and face challenges securing funding for pure code-porting efforts. Given BLAS’ central role and its extensive GPU support, there have been numerous attempts to automate GPU usage, as outlined in Section 2.2. However, limitations inherent to conventional GPU architectures often necessitate frequent data transfers between main memory and GPU memory, resulting in overheads that are unacceptable for practical use.
@@
In our previous work (Li et al., 2024; Li and Wang, 2024), we presented a proof-of-concept framework for symbol interception and replacement using Dynamic Binary Instrumentation (DBI). Preliminary experiments demonstrated the feasibility of automatically offloading *gemm (general matrix multiplication) routines, providing evidence that performant, automatic BLAS offloading is achievable. In this paper, we extend the implementation to support all level-3 BLAS routines and introduce a novel data management strategy, referred to as the Device First-Use policy. Inspired by the OpenMP First-Touch memory management model commonly used in multi-socket CPU or NUMA systems, the Device First-Use policy targets CPU-GPU architectures. It operates by migrating data to device memory (e.g., GPU memory) upon its first access by a GPU kernel. This policy is both conceptually simple and practically effective, significantly reducing redundant data transfers in typical BLAS-intensive workloads. We evaluate the approach across several scientific applications with heavy BLAS usage, scaling up to hundreds of GPU nodes. Results show substantial performance gains. Although our experiments were conducted on the NVIDIA Grace-Hopper system, the proposed methodology is broadly applicable to any CPU-GPU system with cache coherency.
The remainder of this paper is organized as follows: Section 2 reviews the NVIDIA Grace-Hopper unified memory architectures and related BLAS offloading work. Section 3 details the implementation of SCILIB-Accel. In Section 4, we apply the tool to two BLAS-intensive scientific computing codes on up to 200 Grace-Hopper nodes and discuss the results along with performance issues of the NVIDIA Grace-Hopper system. Finally, Section 5 concludes the paper.
@@ 2.1. Coherent memory in NVIDIA Grace-Hopper
Although the two memory subsystems can be accessed cache coherently by both the CPU and GPU, the bandwidth varies significantly between different access patterns, as shown in Table 1. When the CPU accesses its local LPDDR5X memory, it achieves a descent bandwidth of more than 400 GB/s. The GPU, when accessing its local HBM3 memory, delivers substantially higher performance, reaching 3.6 TB/s. The NVLink-C2C interconnect provides 450 GB/s of bandwidth in each direction, which is sufficient to saturate the LPDDR5X bandwidth and allows the GPU to access remote LPDDR5X memory at over 400 GB/s. In contrast, CPU access to HBM3 memory is considerably slower, with measured bandwidth of approximately 140 GB/s. These bandwidth disparities indicate that data locality remains crucial for Grace-Hopper systems. To achieve optimal performance, developers must explicitly manage data placement and movement, similar to conventional GPU architectures, rather than rely solely on coherent unified memory access for production workloads.
@@ 3. Performant Automatic BLAS Offload
Figure 1. Workflow of Automatic BLAS Offload. The BLAS in the CPU binary is intercepted and replaced with a BLAS wrapper where GPU BLAS call is made and data is moved between CPU resident memory and GPU resident memory.
A basic workflow of the automatic BLAS offload tool is illustrated in Figure 1 where a dgemm call in the caller code is intercepted and redirected to a BLAS wrapper that manages data movement and makes the GPU BLAS call. Conceptually, this workflow contains two tasks: 1) intercept BLAS symbols and replace them with a BLAS wrapper where GPU BLAS is called, 2) manage data movement between CPU and GPU resident memories. In the following content, we will discuss how these tasks are implemented in SCILIB-Accel, and most critically how the data movement can be optimized using a novel data movement strategy inspired by the OpenMP First-Touch data placement policy.
@@ 3.1. Symbol Interception
For automatic BLAS offload, we intercept BLAS calls where the trampoline function (BLAS wrapper function) has the same signature as the original function. This design incurs minimal overhead, limited to the cost of two additional jump instructions. This mechanism finds extensive use in profilers, and here we use the PEAK(Wang and Li, 2023) lightweight profiler as the DBI framework, ensuring portability across various architectures including but not limited to x86 and ARM. Inside the DBI framework, the FRIDA-GUM(Frida, nd) binary instrumentation backbend is chosen for simplicity and performance.
@@ 3.2. Data Movement Strategies
Managing data movement is often the most critical part of GPU porting as data transfer speed has been a limiting factor. It is even more so for developing an automatic offload tool, as the tool deals with pure CPU code that is totally untuned for GPU. Here, we discuss three data movement strategies that utilize different features of Grace-Hopper so that this helps us better understand the challenge and opportunities for doing automatic offload on unified memory architecture.
@@ 3.2.1. Strategy 1, Mem-Copy
void my_dgemm(..., host_A, host_B, host_C, ...) {
GPU_Malloc(&device_A,size_A);
GPU_Malloc(&device_B,size_B);
GPU_Malloc(&device_C,size_C);
GPU_Memcpy(host_A to device_A,size_A);
GPU_Memcpy(host_B to device_B,size_B);
GPU_Dgemm(handle, TransA, TransB, m, n, k,
&alpha, device_A, lda, device_B, ldb,
&beta, device_C, ldc);
GPU_Memcpy(device_C to host_C,size_C);
@@ 3.2.2. Strategy 2, counter-based migration
Since the CPU resident memory (LPDDR5X) and the GPU resident memory (HBM3) are physically unified with cache-coherent NVLink C2C, CPU matrix pointers can be passed directly to cuBLAS calls. One can also use numactl to force all memory to be resident on the HBM, but given the poor bandwidth of CPU access HBM, a performance penalty is expected. A new feature in Grace-Hopper designed to better serve the unified memory is the access counter on Hopper GPU, namely the CUDA runtime will automatically move memory pages from LPDDR5X to HBM3 based on remote memory accesses detected by the counter. This counter-based migration mechanism can serve automatic offload when CPU resident matrices are passed to a GPU kernel. An example of a pseudocode is shown below.
@@
void my_dgemm(..., host_A, host_B, host_C, ...) {
GPU_Dgemm(handle, TransA, TransB, m, n, k,
@@ 3.2.3. Strategy 3, Device First-Use policy
One alternative way to look at the Grace-Hopper superchip is that it operates as a heterogeneous dual-socket system, with one socket being a CPU and the other a GPU. Its NUMA configuration also mirrors that of a dual-socket CPU system, where CPU-resident memory is assigned to NUMA 0 and GPU-resident memory to NUMA 1. The data management challenge here in CPU-GPU superchip is a reminiscence of the OpenMP First-Touch data placement policy used in CPU-CPU NUMA programming. While malloc calls can theoretically be intercepted to allocate memory directly for GPU access, these calls originate from the CPU binary, making it impractical to identify which memory regions will later be used by the GPU. As a result, implementing a GPU-first-touch policy is not feasible. To address this, we propose a GPU First-Use policy, where data is moved to the GPU the first time it is accessed by a CUDA kernel. More generally, this concept can be extended to a Device First-Use policy, applicable to any accelerator device with cache-coherent memory access. Table 2 summarizes the features of Device First-Use and highlights its similarities to OpenMP First-Touch.
@@
void my_dgemm(..., host_A, host_B, host_C, ...) {
@@
if (host_C on NUMA 0) move_pages (host_C, to NUMA1) ;
GPU_Dgemm(handle, TransA, TransB, m, n, k,
@@ 3.3. Usage Instructions
    Data management strategies with Device First-Use as default.
@@
    Minimum matrix size to be offloaded.
    If matrices are small, then the BLAS call will stay on CPU. The default threshold is $N_{avg}>500$ where $N_{avg}$ is the average matrix size, the definition of $N_{avg}$ is routine dependent. For general matrix multiplication routines C=AxB, $N_{avg}=(MNK)^{1/3}$ where dimensions of matrices A, B and C are $M\times K$, $K\times N$ and $M\times N$. The default threshold is a safe lower-bound based on preliminary dgemm testings on Grace-Hopper and can certainly be further fine-tuned for different kernels or precisions. The optimal threshold is GPU-dependent. On Grace-Hopper, the default value is a conservative threshold obtained from performance evaluations and guarantees performance gain when running on GPU.
@@ 4. Performance Testings and Discussion
In this section, we perform application tests using production HPC codes. Since quantum chemistry/physics are known to heavily rely on BLAS operations, as linear algebra is the natural language of quantum mechanics, two codes, MuST and PARSEC, from this scientific domain are chosen. These codes are part of the Characteristic Science Application (CSA)(Foundation, 2021) efforts funded by the National Scientific Foundation (NSF) as representative workloads for the Leadership Class Computing Facility.
@@
It is important to emphasize that all performance comparisons presented here are based on an equal number of nodes: Grace-Grace nodes (two CPUs with 144 cores) and Grace-Hopper nodes (one CPU and one GPU). This approach ensures both simplicity and fairness in the analysis. Power and cost are also key factors in performance studies. The power consumption of a Grace-Hopper node is approximately twice that of a Grace-Grace node under full load. Additionally, for each node-hour on the Vista(Texas Advanced Computing Center, 2024) supercomputer, the Service Unit (SU) charged to users for Grace-Hopper nodes is three times higher than that for Grace-Grace nodes. This charge rate reflects the costs associated with acquiring, maintaining, and supporting these different types of nodes.
@@ 4.4. Performance Issues with Grace-Hopper Relevant to Auto Offload
In this part of the paper, we discuss a few performance issues observed in the above application test, these issues reflect the immaturity of the software or hardware design in the current Grace-Hopper system, and could be fixed in the future to further improve performance and usability of automatic offload.
@@ 4.4.1. Counter-based page migration
Table 6. Behavior of CUDA Counter-based Data Migration: by cublasDgemm using host malloc pointers
Matrix Matrix Sizes CPU-¿GPU
Dimensions (in MB) migrated
(M, N, K) A B C A B C
(1000, 1000, 1000) 8 8 8 yes yes yes
(5000, 5000, 5000) 200 200 200 yes? yes? no
(20000, 20000, 20000) 3200 3200 3200 yes no no
(32, 2400, 93536) 24 1796 0.6 yes no no
Result with question mark (?) indicates inconsistent run-to-run behavior.
@@
Since the data migrated to GPU will not be migrated back, automatic offload with counter-based migration should work very similar to manually implemented Device First-Use policy, but we have already seen the slow performance of the counter-based migration in application tests. Here we present a few simple dgemm test cases with different matrix sizes, and provide a deeper understanding of the issues with the counter-based migration. In the following test, matrices A(M,K), B(K,N) and C(M,N) are allocated by malloc and initialized on CPU resident memory, then multiplication $C=A\times B$ is performed at least 5 times by passing these matrices to cublasDgemm, so that the access-counter should allow the matrices to be migrated to GPU resident memory. The NUMA locations of the matrices are reported after each cublasDgemm call, and runtime for each call is reported. From the NUMA location, we can infer whether the data is on the CPU resident memory (NUMA 0) or GPU resident memory (NUMA 1). Four sets of inputs with different matrix sizes and shapes are tested, all of which use FP64 precision.
@@
The case gets strange for non-square matrix dimensions. When M=32, N=2400 and K=93536, a matrix size commonly used in the PARSEC workload, the matrix size becomes 24.0MB for A, 1795.9MB for B, and 0.6MB for C. Throughout our test, only matrix A gets migrated to GPU, while B and C are always on the CPU. This is very counter-intuitive as one would expect the bigger matrix B should at least be moved, as it generates the largest volume of remote memory accesses, but NVIDIA’s counter-based migration algorithm fails to move it.
From these tests, we can see that NVIDIA’s algorithm tends to migrate smaller data to GPU. The reason could be that the algorithm only makes the decision based on a single CUDA kernel call, i.e. moving the big data is costly comparing to that single CUDA kernel runtime, so it decides not to migrate disregarding the fact that the big data gets accessed by the GPU many more times later on. The current counter-based migration is thus unpredictable and inconsistent, and NVIDIA should provide a simple way to disable it by users instead of requiring to unload a kernel module by root.